OpenStudy (anonymous):

Solve the equation and check the solution. -4(5n+1)=8 The solution is n=

OpenStudy (cwrw238):

-20n - 4 = 8 -20n = 8 + 4 = 12 n = 12 / -20 = -0.6

OpenStudy (cwrw238):

check solution -4(5*-0.6) + 1) = -4 * -2 = 8

OpenStudy (anonymous):

Thank you! Looks like I was making it more difficult than it is.

OpenStudy (cwrw238):

yes - its pretty easy really